# Event-Driven Safe and Resilient Control of Automated and Human-Driven Vehicles under EU-FDI Attacks

### 03 Mar 2026 v1

@@ IV-A Performance of the Proposed EDSR Framework under EU-FDI
The simulation results are shown in Figs. 46. Figure 4 demonstrates that CAV $B$ successfully completes the lane-change maneuver within approximately $6.5$ s under EU-FDI attacks, while both CAVs maintain smooth velocity profiles without excessive acceleration or braking. Figure 5 presents the safety constraint values $b_{i,j}$ for all interacting vehicle pairs. All curves remain above zero throughout the maneuver, confirming that the proposed EDSR framework preserves safety and maintains forward invariance of the safe set despite adversarial disturbances. Figure 6 shows the speed tracking errors $\varepsilon_{i}(t)$ of CAVs $A$ and $B$. Although transient deviations appear due to EU-FDI attacks, the errors remain bounded and quickly recover, indicating stable velocity regulation and effective attack mitigation.
@@ IV-B Comparative Study with Event-Driven CBF in [xiao2022event]
Under this unified attack setting, Figure 8 presents representative snapshots of the lane-changing process under the conventional event-driven method at $t=1.50$ s and $t=3.69$ s. The left configuration in Figure 8(a) ($t=1.50$ s) shows that the HDV has already intruded into the safety region of CAV $B$, indicating a violation or near-violation of the safety constraint $b_{B,H}$, while CAV $B$ fails to initiate effective avoidance under EU-FDI attacks. The right configuration in Figure 8(b) ($t=3.69$ s) further reveals that CAV $B$ deviates significantly from the desired lane-changing trajectory, leading to degraded maneuver effectiveness and excessive fallback. These snapshots demonstrate progressive safety degradation in the absence of attack-resilient compensation. Figure 9 shows the lateral positions and speeds of the vehicles. Without resilience to adversarial inputs, small HDV deviations induce drastic maneuvers and severe deceleration in CAV $B$, while CAV $A$ exhibits rapid and uncontrolled acceleration. The disturbance propagates through the traffic interaction, and CAV $B$ loses maneuver feasibility, resulting in premature termination at approximately $t=4.28$ s. This confirms that unmitigated EU-FDI attacks can rapidly drive the system toward unsafe behaviors. Figure 11 presents the CBF safety metrics $b_{i,j}$. The most critical degradation occurs in $b_{B,H}$, which drops below zero between approximately $t=2.15$ s and $t=2.60$ s and reaches about 140-140, clearly indicating loss of forward invariance. Other safety metrics remain comparatively less affected, showing that the CAV BB–HDV interaction dominates the safety degradation. Figure 10 shows the tracking errors $\varepsilon_{i}(t)$. Under escalating EU-FDI attacks, CAV $B$ exhibits severe negative errors (minimum about $30~\mathrm{m/s}$), indicating prolonged excessive deceleration, while CAV $A$ shows sharp positive errors (up to about $15~\mathrm{m/s}$), reflecting unwarranted acceleration. These large error magnitudes confirm that the conventional event-driven method fails to maintain stable speed tracking and produces highly unbalanced vehicle behaviors. Overall, the pure event-driven optimization attempts to preserve safety through extreme acceleration or braking under attacks, which results in impractical trajectories and degraded traffic stability, thereby highlighting the necessity of the integrated resilience provided by the proposed EDSR framework.
@@ V Conclusion
This paper has formulated and addressed the EDSR problem for CAVs interacting with HDVs in mixed traffic environments. To jointly ensure safety and attack resilience under adversarial conditions, a novel control framework has been proposed that integrates event-driven CBFs and CLFs with data-driven estimation of HDV behaviors. This approach maintains provable collision avoidance guarantees during safe lane-changing maneuvers while significantly reducing computational burden in the presence of EU-FDI attacks. Simulation results validate the effectiveness of the framework in ensuring real-time safety and resilient decision-making in uncertain and adversarial traffic scenarios.

### 09 Mar 2026 v2

@@ IV-A Performance of the Proposed EDSR Framework under EU-FDI
The simulation results are shown in Figs. 46. Figure 4 demonstrates that CAV $B$ successfully completes the lane-change maneuver within approximately $6.5$ s under EU-FDI attacks, while both CAVs maintain smooth velocity profiles without excessive acceleration or braking. Figure 5 presents the safety constraint values $b_{i,j}$ for all interacting vehicle pairs. All curves remain above zero throughout the maneuver, confirming that the proposed EDSR framework preserves safety and maintains forward invariance of the safe set despite adversarial disturbances. Figure 6 shows the speed tracking errors $\varepsilon_{i}(t)$ of CAVs $A$ and $B$. The errors remain smooth and bounded throughout the maneuver, indicating stable velocity regulation and effective mitigation of EU-FDI attacks.
@@ IV-B Comparative Study with Event-Driven CBF in [xiao2022event]
Under this unified attack setting, Figure 8 presents representative snapshots of the lane-changing process under the conventional event-driven method at $t=1.50$ s and $t=3.69$ s. The left configuration in Figure 8(a) ($t=1.50$ s) shows that the HDV has already intruded into the safety region of CAV $B$, indicating a violation of the safety constraint $b_{B,H}$, while CAV $B$ fails to initiate effective avoidance under EU-FDI attacks. The right configuration in Figure 8(b) ($t=3.69$ s) further reveals that CAV $B$ deviates significantly from the desired lane-changing trajectory, leading to degraded maneuver effectiveness and excessive fallback. These snapshots demonstrate progressive safety degradation in the absence of attack-resilient compensation. Figure 9 shows the lateral positions and speeds of the vehicles. Without resilience to adversarial inputs, small HDV deviations induce drastic maneuvers and severe deceleration in CAV $B$, while CAV $A$ exhibits rapid and uncontrolled acceleration. The disturbance propagates through the traffic interaction, and CAV $B$ loses maneuver feasibility, resulting in premature termination at approximately $t=4.28$ s. This confirms that unmitigated EU-FDI attacks can rapidly drive the system toward unsafe behaviors. Figure 10 shows the tracking errors $\varepsilon_{i}(t)$. Under escalating EU-FDI attacks, CAV $A$ shows sharp positive errors (up to about $15~\mathrm{m/s}$), reflecting unwarranted acceleration, while CAV $B$ exhibits severe negative errors (minimum about $30~\mathrm{m/s}$), indicating prolonged excessive deceleration. These large error magnitudes confirm that the conventional event-driven method fails to maintain stable speed tracking and produces highly unbalanced vehicle behaviors. Figure 11 presents the CBF safety metrics $b_{i,j}$. The most critical degradation occurs in $b_{B,H}$, which drops below zero between approximately $t=2.15$ s and $t=2.60$ s and reaches about 140-140, clearly indicating loss of forward invariance. Overall, the conventional event-driven CBFs method attempts to preserve safety through extreme acceleration or braking under attacks, which results in impractical trajectories and degraded traffic stability, thereby highlighting the necessity of the integrated resilience provided by the proposed EDSR framework.
@@ V Conclusion
This paper formulates and addresses the SAR control problem for CAVs interacting with HDVs in mixed traffic environments. To jointly ensure safety and attack resilience under adversarial conditions, a novel EDSR control framework has been proposed that integrates event-driven CBFs and CLFs with data-driven estimation of HDV behaviors. This approach maintains provable collision avoidance guarantees for safe lane-changing maneuvers even in the presence of EU-FDI attacks, while also significantly reducing the computational burden through the event-driven design. Simulation results validate the effectiveness of the framework in ensuring real-time safety and attack-resilient control in uncertain and adversarial traffic scenarios.